OpenStudy (anonymous):

write the partial fraction decomposition of the rational expression. Check your result algebraically. 8/ x^2-4x-21

OpenStudy (tkhunny):

You already have it. \(\dfrac{8}{x^{2}} - 4x - 21\) Of course, if you meant \(\dfrac{8}{x^{2} - 4x - 21}\), which you did NOT write, you should first factor the denominator.

OpenStudy (anonymous):

oh, sorry/ I meant the second one you had wrote.

OpenStudy (tkhunny):

Sweet. Remember your Order of Operations. Use parentheses as necessary to clarify intent. Factor away!

OpenStudy (anonymous):

did you get \[8=A(x-7)+B(x+3)\]?

OpenStudy (anonymous):

i can show you a real snappy way to do this if you like

OpenStudy (anonymous):

start with \[\frac{8}{(x-7)(x+3)}=\frac{A}{x-7}+\frac{B}{x+3}\] now lets find \(A\) quickly

OpenStudy (anonymous):

Sure! I would like to find shortcuts! ^_^

OpenStudy (anonymous):

the denominator is of \(A\) is \(x-7\) and that would be zero if \(x=7\) so on the left we are going to cover up the factor of \(x-7\) with your finger, and replace \(x\) by \(7\) i.e. \[\frac{8}{\cancel{x-7}(7+3)}=\frac{8}{10}=\frac{4}{5}=A\]

OpenStudy (anonymous):

similarly to find \(B\) the denominator is \(x+3\) which is zero if \(x=-3\) so cover up that factor on the left, replace \(x\) by \(-3\) and get \[\frac{8}{(-3+7)\cancel{x+3}}=\frac{8}{-10}=-\frac{4}{5}=B\]

OpenStudy (anonymous):

i made a typo there, should have been \[\frac{8}{(-3-7)\cancel{x+3}}=\frac{8}{-10}=-\frac{4}{5}=B\]

OpenStudy (anonymous):

Thank you so much! You are a helpful and brilliant person.

OpenStudy (anonymous):

lol yw